OpenStudy (anonymous):

Find the partial decomposition of -3x^2 + 13x -12 all over x^3 - 4x^2 +4x

OpenStudy (anonymous):

im thinking you have to factor out the denominator?

OpenStudy (anonymous):

factor out an x in x^3 - 4x^2 +4x

OpenStudy (zarkon):

\[x^3-4x^2+4x=x(x-2)^2\]

OpenStudy (anonymous):

(-3x^2 + 13x -12)/(x(x-2)^2)

OpenStudy (anonymous):

yea.. i forgot how to factor cube

OpenStudy (anonymous):

A/x + B/(x-2) + C/(x-2)^2=(-3x^2 + 13x -12)/(x(x-2)^2)

OpenStudy (anonymous):

whats the next step?

OpenStudy (anonymous):

what do you think is the next step?

OpenStudy (anonymous):

i wrote A(x-2)(x-2)^2 + Bx(x-2)^2 + Cx(x-2) = -3x^2 +13x -12

OpenStudy (anonymous):

i think its wrong

OpenStudy (anonymous):

Multiply both sides by (x (x-2)²)

OpenStudy (anonymous):

looks good to me except for the \[Bx(x-2)^2\] part. think that should be \[Bx(x-2) \] yes?

OpenStudy (anonymous):

\[A(x-2)^2 + Bx(x-2) + Cx = -3x^2 +13x -12 \]

OpenStudy (anonymous):

right, now let x=0 and solve for A

OpenStudy (anonymous):

\[A(-2)^2=-12\] \[4A=-12\] \[A=-3\] how'd i do?

OpenStudy (anonymous):

that looks gooD!

OpenStudy (anonymous):

now guess what you let x =?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

how about b? do i plug in for A and C?

OpenStudy (anonymous):

did you find C?

OpenStudy (anonymous):

C = 1

OpenStudy (anonymous):

think b = 0

OpenStudy (anonymous):

ok so now you have \[-3(x-2)^2+Bx(x-2)+x=-3x^2+13x-12\]

OpenStudy (anonymous):

yes B = 0

OpenStudy (anonymous):

can you help me with another problem im going to make a new post

OpenStudy (anonymous):

sure make a new one

OpenStudy (anonymous):

btw answer is \[\frac{1}{(x-2)^2}-\frac{3}{x}\]

OpenStudy (anonymous):

ok